OpenStudy (anonymous):

11. In an 8 inch square cake pan and a 9 inch square cake pan, what is the difference in volume each pan will hold? Assume each pan is 3 inches high.

OpenStudy (richyw):

volume of a rectangular prism is \[V=lwh\]where l=length, w=width and h=height. This is a square pan so \(l=w\). Difference in volume is given by\[\Delta V=|V_2-V_1|\]

OpenStudy (richyw):

so calculate each volume (\(V_1\) and \(V_2\) ) using the first formula. Then calculate the difference using the second formula

OpenStudy (anonymous):

lwh=9*8*3=216 then what is next

OpenStudy (richyw):

that is incorrect. Remeber you are calculating the volume for TWO pans. Both have square bottoms, one is a 9x9 square, the other is a 8x8 square.

OpenStudy (richyw):

so for example your first volume will be \(V_1=9*9*3=243\)

OpenStudy (richyw):

now calculate \(V_2\) the volume of the 8 by 8 inch pan...

OpenStudy (anonymous):

8*8*3=192

OpenStudy (anonymous):

then do you subtract the two 243-192=51